OpenStudy (anonymous):

solve for y. -1 + y 8 < 15 y < 7 4 y < 2 y < 112 y < 128

OpenStudy (anonymous):

\[8<\frac{ y }{ 8 }\] \[y <2\] \[y <112\] \[y <128\]

OpenStudy (anonymous):

Your question is 8<y/8?

OpenStudy (anonymous):

Multiply both sides by 8

OpenStudy (anonymous):

no the question is\[-1+\frac{ 7 }{ 8 }<15\]

OpenStudy (anonymous):

heh

OpenStudy (anonymous):

the first comment are the choices.

OpenStudy (anonymous):

Is it 7/8y?

OpenStudy (anonymous):

\[-1+\frac{ 7 }{ 8 }y < 15\]

OpenStudy (anonymous):